OpenStudy (anonymous):

How do I find the rate of an exponential function (with a graph) over a given interval?

OpenStudy (amistre64):

y = B r^(x)

OpenStudy (amistre64):

y/B = r^x (y/B)^(1/x) = r

OpenStudy (anonymous):

hey

OpenStudy (anonymous):

Oh pellet.. My mind just got blownnn.... Sorry but math isn't my strongest subject.

OpenStudy (anonymous):

Mind blown?! Is you okay?!

OpenStudy (anonymous):

I is good! :)

OpenStudy (anonymous):

I would rather not have to read cuss words please.

OpenStudy (anonymous):

My bad...

OpenStudy (amistre64):

if we assume y = 3(2)^x (1,6) and (2,12) are two points that we can work with to see if we can get back to that setup (6/B)^(1/1) = r (12/B)^(1/2) = r 6/B = (12/B)^(1/2) 36/B^2 = 12/B 36/12 = B^2/B 3=B (6/3)^(1/1) = r = 2

OpenStudy (anonymous):

We can still be friends.

OpenStudy (anonymous):

Lol oookkkkk.

OpenStudy (amistre64):

can you define 2 points on the graph?

OpenStudy (anonymous):

can i be your friend @Ben_Henderson ??????? lol

OpenStudy (anonymous):

I don't have any practice problems @amistre64

OpenStudy (anonymous):

yahhhhh @armine

OpenStudy (anonymous):

Do yall have an instagram? @armine @niallhoranbabe<3

OpenStudy (amistre64):

your post mentions that you have a graph to work with

OpenStudy (anonymous):

Yess! Whats yours? I'll add you

OpenStudy (anonymous):

yeah lol mine is smile4zoe2 i think lol

OpenStudy (anonymous):

I'm private so I'll have you accept your request.

OpenStudy (anonymous):

Ok!

OpenStudy (anonymous):

BRB

OpenStudy (anonymous):

Lol okay. Im aadding u @armine

OpenStudy (anonymous):

k

OpenStudy (anonymous):

I'm gonna log off. I will probably log back on at like 6 or 7. BYEEEEEE

OpenStudy (anonymous):

Eastern Time

OpenStudy (anonymous):

BYE! I added u on instagram btw!

OpenStudy (anonymous):

Thanks! You is okay! LOLOL

OpenStudy (anonymous):

I IS GOOOODDD!!!

OpenStudy (anonymous):

Byee